OpenStudy (anonymous):

Simplify the given expression. square root -10 * square root -8

OpenStudy (anonymous):

Is this you expresion? \[\sqrt{-10}\cdot\sqrt{-8}\]?

OpenStudy (anonymous):

Remember that \[\sqrt{-1}=i

OpenStudy (mertsj):